Prev Next

SW&IT Meeting Minutes 2021-04-05

Attendees: David, Matthieu, Michael, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20210308
LR20210308 - Bec, Matthieu 80.0
LR20210308 - Edmeades, David 82.0
LR20210308 - Sola, Igor 0.0
LR20210308 - Sizemore, Michael 13.0
LR20210308 - Grenz, Paul 40.0
LR20210308 - Hooper, Stephen 80.0
LR20210308 - Zhang, Yang 70.0
LR20210308 - Zhang, Xianyu 64.0

News from management

  • Value of the week to discuss: Integrity - Our words and our actions are a direct expression of who we are, and through them we model our core values.

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management
    • ATT WG, PM FTE & priorities
    • secured websockets
    • OT SHARK-NIR
  • This week
    • candidates tech questionaire and follow up interviews
    • websocket demo and wrappers for publishers
    • sci-arch meeting

DE

  • Last week
    • Synology
      • Several false starts, finally was able to get the update error sorted and the replacement unit installed
    • Cisco training
    • Storage usage info compile
    • Cloud backup work
    • Emailed FM about blower power/direction in server room
  • This week
    • Synology return
      • SH still to go in today? If not, I have a visit okayed.
    • Archive meeting Tues
    • Big Cisco push
    • More resource usage info compiling
    • Find a time to collect Igor's hardware

YZ

  • Previous:
    • OSCO dms page officially released (many fixes/tweaks/new features)!
    • TMS work on extraction utility upgrade, data analysis, wiki work.
    • Some RB_Science trouble-shooting.
    • Laura's ATT/AO survey.
    • Did some safety training.
  • Next:
    • TMS work:
      • Start database transition work.
      • Drafting TMS PC software redesign requests.
      • Some wiki work.
    • OSCO work:
      • DMS anouncement.
      • OSCO background process tweaking.
      • Diagram for Al.
    • RB_Science patch again.
    • MODS 64-bit stuff.
    • More safety training.

XZ

  • Last week
    • ATT AdSec /SW
    • SX ASM operation support
    • DX ASM work preparation
    • SOUL SW support
    • AO support
  • This week
    • SX ASM operation issues
    • DX ASM work preparation (ask IT for remote power switch)
    • Safety training
    • AO support

PSG

  • Last week
    • Got COVID shot, ended up taking sick days and couldn't perform tasks.
  • This week
    • Coordinate with Dan Cox regarding flash drive backup.
    • See if spare controller unit has arrived.
    • Fix UMAC controller code.
    • Arrange a time to be in the lab with Juan Pable to test said code.

SH

  • DIMM with LeRoy, broken fiber
  • Some UA Secure
  • Docker upgrade to F33
  • SSL certificates, rearrangement and LetsEncrypt
  • abell upgrade

next:

  • UITS stuff
  • PEPSI VM
  • some more VM stuff
  • DIMM with LeRoy (maybe)
  • LDAP
Topic revision: r9 - 05 Apr 2021, StephenHooper
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback